THE MEASURE OF AN ANGLE IS 28 DEGREES GREATER THAN IT'S COMPLEMENT. FIND THE MEASURE OF EACH ANGLE.
THE MEASURE OF AN ANGLE IS 21 DEGREES MORE THAN TWICE IT'S SUPPLEMENT. FIND THE MEASURE OF EACH ANGLE.

Answer by rfer(16322)

x+x+28=90
2x=90-28
2x=62
x=31
x+28=59
-------------
x+2x+21=180
3x=180-21
3x=159
x=159/3
x=53
2x+21=127